# Examining the Utility of Self-disclosure Types for Modeling Annotators of Social Norms

### v1

@@ 6 Analysis and Discussion
For RQ1, we hypothesized that a larger number of comments would work best and that sentences would be more focused than comments. Similar samples did not significantly outperform random samples, though there were slight differences. However, in contrast to previous work, a smaller number of comments were needed. It appears that sentences, while possibly more focused, may be more highly variable, whereas comments provide a more substantial amount of information. It may also be the case that generation requires more examples to mimic the style, whereas fewer samples are needed to predict verdicts, given that those samples are related. We further examined the correlation between the length of comments in tokens and accuracy for 20 similar comments (highest F1 model in Table 4), finding a very weak Pearson $r=0.01$ but significant ($p<0.04$) correlation, suggesting the comment length has little to do with how accurately verdicts are classified.
@@
As mentioned in §5, even a large number of sentences still underperforms a small number of similar posts. If we consider similar sentences to be only the unit of information that contains a fact about the annotator, it is more similar to work on personalization that uses personal attributes, as discussed in §2. This type of metadata contains an age or gender, whereas the longer posts from annotators may contain rich information that perspectivist classifiers can leverage. While it is not strictly the case that sentences from our corpus represent isolated facts, this reasoning supports, and may partially explain why posts are more useful than sentences.
@@
Next, we found that our ranking of theory-based categories was correct (RQ2). Demographics was the highest performing category followed by attitudes, though experiences and relationships tied for lowest F1. Lastly, for RQ3, we found that some automatic categories performed almost as well as theory-based categories, but that theory-based approaches (mostly demographics and attitudes) outperformed clusters. It was not as intuitive to understand why one cluster outperforms another, e.g. why relationship issues with men performed much better than issues with women. Similar amounts of effort went into constructing both theory-based and automatic categories. Future work will need to run more extensive experiments to find more comprehensive answers to these questions, but it appears that defining self-disclosure types using theory can lead to improved performance.
@@
Sampling Analysis. We further examine the type of similar posts used to represent the annotators. For our experiments with all comments, we plot the proportion of times that the five most similar posts used to represent that annotator come from each category, shown in Figure 2. We note that 75% of these top-5 similar posts have no theory-based category and 78% of them have no automatic category. Using either of our methods to categorize these posts eliminates the majority of similar comments. This explains the slightly lower performance when using comments from only one category rather than sampling from the full pool of posts from that author. We include an additional analysis of subcategories in Appendix C.
@@
We further examined the five most similar posts to find out how diverse the set of sampled posts was and the imbalance in the distribution of the most frequently retrieved posts. We include Figure 3 in Appendix D, which shows that a median of 33% of an authors posts end up in the set of similar posts at some point. Furthermore, the ratio of the most frequent similar post to the second most frequent has a median of 1.2. This suggests that a diverse set of comments are used to represent the annotators across situations.
@@
We further observed that demographics were the highest performing category. This group contains identities, which use the most generic regular expressions to match self-disclosures, making the group more diverse. The experiments from Table 4 show higher performance than when using any individual type of information by up to 4.2 points F1. Further, Figure 2 shows that a variety of posts are sampled in these experiments. Diverse sampling significantly outperforms the all posts method from previous work. This evidence supports that a diverse sample of information from an individual greatly improves performance.
@@ 7 Conclusions
We explored experimental settings for predicting verdicts based on self-disclosure statements. We found that contrary to prior work, only a small number of similar comments are needed to effectively model annotators and that in contrast, using many or all posts from an annotator performed worse. When examining the type of information, we found that theory-based categories performed similarly to automatic categories (clusters), but that theory-based categories were more consistently high performing. Demographics outperformed all other types of self-disclosures. Our analysis suggests that it is much more important to have a diverse sample of different types of annotator information. Our work suggests entirely new avenues of research for improving annotator modeling outside of what is normally collected during annotation. As posts also outperform sentences, even when a large number of sentences are used, we suggest that the unstructured nature of our texts may provide important information in latent factors for representing users and suggest that future work further explores annotator writing. We will release our code and data to support future work on annotator modeling.
@@
Self-disclosure statements were extracted with regular expressions that will not catch all cases of statements of a given type. They will also capture false positives. There are ways to improve the rule-based extraction of self-disclosures, though we found ours had a low enough false-positive rate to be useful for our experiments. One example of this is with the category of possessions, where one can say “I have X,” but X could be “had concerns,” or otherwise describe the past rather than a present object or experience. Rules with parsing features could make this more specific.
We did not account for the change in attributes over time. If a poster has posted over many years, their age or gender identity would more likely change leading to a set of conflicting statements. These types of issues could be resolved with rules, e.g. by taking the most recent values, however the most recent values may not correspond to their thoughts when providing a verdict. One could argue that a bag of identity statements is more representative of an individual (i.e. they were active posting at a range of ages, gender identities), however, we are unaware of the magnitude of the impact of rule-based filtering of these statements.
It is difficult to account for all factors in a fully controlled experiment due to the interdependent nature of this kind of real world data. The number of annotators with at least five posts belonging to all automatic and theory-based categories was only 537 or 2.69% of the total author set. For this reason, we cannot use the same number and type of available self-disclosures for each condition. When an annotator has less than five, we use the full available set and there is effectively no impact of the similarity function, as there are no posts to selectively exclude. These criteria may help future researchers in designing new data collection protocols.
@@
Any work that attempts to model individual’s their opinions, language patterns, or behaviors has the potential to be misused. Modeling individual views could be used adversarially to find ways to present conflict situations that do not incite opposition, or conversely, how to present conflicts that maximizes opposition for a target individual or group.
@@ Appendix B Self-disclosure Exploration
In this appendix, we discuss additional explorations of the self-disclosure statements. We also analyzed the most common unigrams, bigrams, and trigrams around self-disclosure phrases, grouped authors by post count, tagged parts of speech to identify frequent patterns and verbs. When examining unigrams, bigrams, and trigrams, we looked at every word in a sentence before and after a self-disclosure phrase and identified the most common unigram, bigram, and trigrams for each. We found that the most common unigram before a self-disclosure is ’i’ (Fig. 8), and is ’to’ after (Fig. 9). The most common bigram before a self-disclosure is "i don’t" (Fig. 10), and is ’to be’ after (Fig. 11). The most common trigram before a self-disclosure phrase is "i don’t think" (Fig. 12), and is ’a lot of’ after (Fig. 13). The three most common Parts of Speech are NN, PRP, and IN (Fig. 6). The three most common verbs are ’is’, ’have’, and ’be’ (Fig. 7). The most common self-disclosure phrases were "I’m", "I have", and "I am". The most common verbs are "was", "is", and "have". The most common parts of speech are nouns, personal pronouns, and Preposition or subordinating conjunction. Additionally, we used PCA to reduce the dimension of our SBERT embeddings, from 384 (SBERT embedding dimension) to two, to plot the kmeans clusters  5. The PCA variance ratio for the first dimension is 0.043 and for the second dimension is 0.034.
33.6020204040606080801001001.20.50.5111.51.522
050100AgeGenderIdentity050100HobbyPossessionWork
@@
Appendix C Analysis of Subcategories and Similar Comments
In our experiments with each category, we also looked at the proportion of times that posts were sampled from the low-level theory-based categories. These are shown in Figure 4. We see that similar posts most often select identity statements, rather than gender and age when using demographics. When we look at experiences, we see that most similar posts come from the set of possessions, with a lower proportion coming from work and hobbies. These proportions roughly follow the frequency of these types of posts.
Figure 3 shows the box plot for the proportion of comments covered in 5 similar comment experiments and the relative frequency of the most frequent similar comment to the second most frequent similar comment as discussed in §6.
Appendix D Cluster Silhouette Scores
The automatic clusters we used were computed using k-means with ten clusters. We reduced the dimensionality using truncated singular value decomposition to reduce the dimension from 384 to 5. After reduction, we compute silhouette scores and found an average score of 0.25. While this is not as high as conventional wisdom suggests is a good score, it serves as evidence that the data is clustered to some extent. Our clusters do overlap to some degree, but most importantly, we also find that they show strong performance on our downstream task of predicting verdicts. The plots for silhouettes are shown in Figure 15.
Figure 15: Silhouette scores for our automatic clusters.
Appendix E Other Splits
@@
Verdict split. This split randomly assigns verdicts to the training, validation, and test sets. While straightforward, it introduces two potential sources of overlap: the same situations and the same annotators may appear in multiple splits. Our dataset forms a fully connected bipartite graph between posts and annotators, where each edge represents a comment. Because many annotators comment on many posts, it’s not possible to remove both sources of overlap at the same time without disconnecting the graph or discarding large portions of data. This split is equivalent to the "no disjoint" setting described in Plepi et al. (2022).
@@
Appendix F Discussion of Automatic Cluster Labels
@@ F.1 C1: Financial Issues
The ’Financial Issues’ cluster contains comments which discuss topics like money, budgets, loans, or directly mention the cost of something, for example \$20.
@@
Centeroid Comment: "NTA. I totally get both sides. I feel uncomfortable when people spend too much money on me too. It’s a pride thing I guess. You work your tail off and still aren’t in a solid financial position. … I know where I’m at, \$1,200 is around what 2.5 weeks of rent + utilities for a family of 5 would cost anyway, you’re just paying it to her instead of a landlord."
@@
Random Comment: "NTA - please DO NOT sell that car! I think you’ll regret it if you do. You love your sister and want to help, but it’s her choice to study abroad and she needs to find a way to make it happen. Why should you have to sacrifice something so special just because you share dna and she wants money."
F.2 C2: Accommodating Social Needs
@@
The ’Accommodating Social Needs’ cluster contains comments which discuss topics centered around Accommodating others, social obligation, and social norms.
@@
Centeroid Comment: "YTA. I’m like this. Partner of many years had never been annoyed by it (he even helps me keep track of where the nearest bathrooms are and asks if I need to go before we leave)"
@@
Random Comment: "I’m glad I could help. I hope you find healing."
F.3 C3: Parenting & Discipline
@@
The ’Parenting & Discipline’ cluster contains comments which discuss topics like general parenting guidelines/styles, disciplining children, and advice for parents.
@@
Centeroid Comment: "I think you may be downplaying how much a parent can ignore or not see their own child’s bad behavior. I have a feeling that’s what is going on here."
@@
Random Comment: "… I have literally NEVER witnessed any new parents doing something so selfish and self-absorbed. Like I said, and like you said, it’s THEIR baby and they have every right to do whatever they want. But it WILL damage their relationships with the people they deem ’close’ to them."
F.4 C4: Family Structure Change
@@
The ’Family Structure Change’ cluster contains comments which discuss topics mainly centered around changes in the family, either adding or subtracting in some way. Examples of this include moving away, divorce, death, or birth in the family which changes how the family operates.
@@
Centeroid Comment: "Love how OP tries to frame it as "looking out for his son" when in reality, he selfishly wants his son to be a doormat & do what he wants versus being happy. I’m betting the daughter got the same treatment & that’s why she went NC to avoid all his stupid drama. … They’ve grown up & learned to be adults, about damn time you do the same before you torpedo any chances of ever having them in your lives again. …"
@@
Random Comment: "Husband has been clear about not wanting another baby but OP has insisted on carrying to term. I think he’s been clear about not wanting to parent."
F.5 C5: Relationship Issues with Men
@@
The ’Relationship Issues with Men’ cluster contains comments which discuss situations in which people have problems with men.
@@
Centeroid Comment: "I think you and him need to have a serious conversation about this. That behavior is unacceptable."
@@
Random Comment: "… In short, you should be able to watch your CNN at his house and he should be able to watch his Fox News. This has nothing to do with his cancer diagnosis, nor does it have to do specifically with your relationship to him. …"
F.6 C6: Relationship Issues with Women
@@
The ’Relationship Issues with Women’ cluster contains comments which discuss situations in which people have problems with women.
@@
Centeroid Comment: "NTA that sounds a little mean to be honest. I think it’s understandable to be upset by it, but hopefully she will take it seriously and stop doing that."
@@
Random Comment: "… I would NEVER talk to my boyfriend that way for trying to help me. You were right to leave. You were being an excellent boyfriend, but she was being a jerk."
F.7 C7: Sympathy & Support

### v2

@@ 6 Analysis and Discussion
For RQ1, we hypothesized that a larger number of comments would work best and that sentences would be for more focused than comments. We found that similar samples significantly outperformed random samples. In contrast to previous work, a smaller number of comments were needed. It appears that sentences, while possibly more focused, may be more highly variable, whereas comments provide a more substantial amount of information. It may also be the case that generation, as performed in Plepi et al. (2024), requires more examples to mimic style, whereas fewer samples are needed to predict verdicts, given that those samples are related. We examined the correlation between the length of comments in tokens and accuracy for 20 similar comments, finding a very weak Pearson $r=0.01$ but significant ($p<0.04$) correlation, suggesting the comment length has little to do with how accurately verdicts are classified.
@@
As mentioned in §5, even a large number of sentences still underperforms a small number of similar posts. If we consider similar sentences to be only the unit of information that contains a fact about the annotator, it is more similar to work on personalization that uses personal attributes, as discussed in §2. This type of metadata contains an age or gender, whereas the longer posts from annotators may contain rich latent information that perspectivist classifiers can leverage. While it is not strictly the case that sentences from our corpus represent isolated facts, this reasoning supports, and may partially explain why comments are more useful than sentences. It also helps address the issue that sentences by themselves, are not always self-contained, as mentioned in §3.2.
@@
Next, we found that our ranking of theory-based categories was correct (RQ2). Demographics was the highest performing category followed by attitudes, experiences, and lastly relationships. Lastly, for RQ3, we found that some automatic categories performed as well as theory-based categories, but that theory-based approaches (mostly demographics and attitudes) outperformed most clusters. It was not as intuitive to understand why one cluster outperforms another, e.g. why parenting and discipline (C3) outperforms relationship issues (C5 & C6). Similar amounts of effort went into constructing both theory-based and automatic categories. Future work will need to run more extensive experiments to find more comprehensive answers to these questions, but it appears that defining self-disclosure types using both approaches can lead to improved performance.
@@
Sampling Analysis. We further examine the type of similar posts used to represent the annotators. For our experiments using all annotator comments, we plot the proportion of times that the five most similar comments used to represent that annotator come from each category, shown in Figure 2. We note that 75% of these top-5 similar comments have no theory-based category and 78% of them have no automatic category. Using either of our methods to categorize these posts eliminates the majority of similar comments. This explains the slightly lower performance when using comments from only one category rather than sampling from the full pool of posts from that author (as in Table 4.
We also looked at the proportion of times that posts were sampled from the low-level theory-based categories. These are shown in Figure 4. We see that similar posts most often select identity statements, rather than gender and age when using demographics. When we look at experiences, we see that most similar posts come from the set of possessions, with a lower proportion coming from work and hobbies. These proportions roughly follow the frequency of these types of posts.
@@
We further examined the five most similar comments to assess the diversity of the sampled comments and the degree of imbalance in the distribution of the most frequently retrieved comments. Figure 3 shows the relative frequency of the most frequent similar comment to the second most frequent similar comment with a median ratio of 1.2. This suggests that a diverse set of comments are used to represent the annotators across different situations. We also see a median of 33% of an authors comments end up in the set of similar comments at some point.
@@
We further observed that the theory-based category, demographics, was the highest performing category. This group contains identities, which use the most generic regular expressions to match self-disclosures, making the group more diverse. Experiments from Table 4 show that diverse sampling with theory-based categories is often outperformed by individual categories, both theory-based and automatic. As shown in Figure 2, we see that a variety of comments are sampled in these diverse experiments. Diverse sampling significantly outperforms the all posts method from previous work, but not the method that uses similar comments from the full set. This evidence suggests that a diverse sample of information from an individual may improve performance, but also that we may still not be capturing the most useful information about individuals, as the similar comment set is much larger than what we are able to extract with either categorization approach.
@@ 7 Conclusions
We explored experimental settings for predicting verdicts based on self-disclosure statements. We found that contrary to prior work, only a small number of similar comments are needed to effectively model annotators and that in contrast, using many or all posts from an annotator performed worse. When examining the types of information that are most beneficial, we found that theory-based categories performed similarly to automatic categories (clusters), but that theory-based categories were more consistently high performing. Demographics outperformed all other types of self-disclosures. While our categories showed notable differences in performance, they all underperformed the five similar comments sampled from the full pool of annotator comments, suggesting that we are still not capturing the most useful information predicting verdicts. The Reddit corpus allows us to explore the use of information that is not normally collected in the annotation process. We find that modeling certain kinds of information can be more informative when modeling individual annotators, suggesting that collecting a wider array of information about annotators during the annotation process could lead to ability to develop improved models. We release our code and data to support future work on annotator modeling. [^1] https://github.com/KieranHenderson/Self-disclosuresForAnnotatorModeling
@@
Self-disclosure statements were extracted with regular expressions that will not catch all cases of statements of a given type. They will also capture false positives. There are ways to improve the rule-based extraction of self-disclosures, though we found ours had a low enough false-positive rate to be useful for our experiments. One example of a false positive is with the category of possessions, which captures the pattern “I have X” but X could be “had concerns,” or otherwise describe the past rather than a present object or experience. Rules with parsing features could make this more specific.
We did not account for the change in attributes over time. If a poster has posted over many years, their age or gender identity would be more likely to change leading to a set of conflicting statements. These types of issues could be resolved with rules, e.g. by taking the most recent values, however the most recent values may not correspond to their thoughts when providing a verdict. One could argue that a bag of identity statements is more representative of an individual (i.e. they were active posting at a range of ages, gender identities), however, we are unaware of the magnitude of the impact of rule-based filtering of these statements.
It is difficult to account for all factors in a fully controlled experiment due to the interdependent nature of this kind of real world data. The number of annotators with at least five posts belonging to all automatic and theory-based categories was only 537 or 2.69% of the total author set. For this reason, we cannot use the same number and type of available self-disclosures for each condition. When an annotator has less than five comments in a given category, we use the full available set and the similarity function has effectively no impact, as there are no posts to selectively exclude. These criteria may help future researchers in designing new data collection protocols.
@@
Any work that attempts to model individual’s, their opinions, language patterns, or behaviors has the potential to be misused. Modeling individual views could be used adversarially to find ways to present conflict situations that do not incite opposition, or conversely, how to present conflict that maximizes opposition for a target individual or group.
@@ Appendix B Self-disclosure Exploration
In this appendix, we discuss additional explorations of the self-disclosure statements. We also analyzed the most common unigrams, bigrams, and trigrams around self-disclosure phrases, grouped authors by post count, tagged parts of speech to identify frequent patterns and verbs. When examining unigrams, bigrams, and trigrams, we looked at every word in a sentence before and after a self-disclosure phrase and identified the most common unigram, bigram, and trigrams for each. We found that the most common unigram before a self-disclosure is “i” (Fig. 8), and is “to” after (Fig. 9). The most common bigram before a self-disclosure is “i don’t” (Fig. 10), and is “to be” after (Fig. 11). The most common trigram before a self-disclosure phrase is “i don’t think” (Fig. 12), and is “a lot of” after (Fig. 13). The three most common Parts of Speech are NN, PRP, and IN (Fig. 6). The three most common verbs are “is”, “have”, and “be” (Fig. 7). The most common self-disclosure phrases were “I’m,” “I have”, and “I am”. The most common verbs are “was”, “is”, and “have”. The most common parts of speech are nouns, personal pronouns, and Preposition or subordinating conjunction. Additionally, we used PCA to reduce the dimension of our SBERT embeddings, from 384 (SBERT embedding dimension) to two, to plot the kmeans clusters  5. The PCA variance ratio for the first dimension is 0.043 and for the second dimension is 0.034.
@@
Appendix C Other Splits
@@
Verdict split. This split randomly assigns verdicts to the training, validation, and test sets. While straightforward, it introduces two potential sources of overlap: the same situations and the same annotators may appear in multiple splits. Our dataset forms a fully connected bipartite graph between posts and annotators, where each edge represents a comment. Because many annotators comment on many posts, it’s not possible to remove both sources of overlap at the same time without disconnecting the graph or discarding large portions of data. This split is equivalent to the “no disjoint” setting described in Plepi et al. (2022).
@@
Appendix D Discussion of Automatic Cluster Labels
@@ D.1 C1: Financial Issues
The Financial Issues cluster contains comments which discuss topics like money, budgets, loans, or directly mention the cost of something, for example \$20.
@@
Centeroid Comment: “NTA. I totally get both sides. I feel uncomfortable when people spend too much money on me too. It’s a pride thing I guess. You work your tail off and still aren’t in a solid financial position. … I know where I’m at, \$1,200 is around what 2.5 weeks of rent + utilities for a family of 5 would cost anyway, you’re just paying it to her instead of a landlord.”
@@
Random Comment: “NTA - please DO NOT sell that car! I think you’ll regret it if you do. You love your sister and want to help, but it’s her choice to study abroad and she needs to find a way to make it happen. Why should you have to sacrifice something so special just because you share dna and she wants money.”
D.2 C2: Accommodating Social Needs
@@
The Accommodating Social Needs cluster contains comments which discuss topics centered around Accommodating others, social obligation, and social norms.
@@
Centeroid Comment: “YTA. I’m like this. Partner of many years had never been annoyed by it (he even helps me keep track of where the nearest bathrooms are and asks if I need to go before we leave)”
@@
Random Comment: “I’m glad I could help. I hope you find healing.”
D.3 C3: Parenting & Discipline
@@
The Parenting & Discipline cluster contains comments which discuss topics like general parenting guidelines/styles, disciplining children, and advice for parents.
@@
Centeroid Comment: “I think you may be downplaying how much a parent can ignore or not see their own child’s bad behavior. I have a feeling that’s what is going on here.”
@@
Random Comment: “… I have literally NEVER witnessed any new parents doing something so selfish and self-absorbed. Like I said, and like you said, it’s THEIR baby and they have every right to do whatever they want. But it WILL damage their relationships with the people they deem ‘close’ to them.”
D.4 C4: Family Structure Change
@@
The Family Structure Change cluster contains comments which discuss topics mainly centered around changes in the family, either adding or subtracting in some way. Examples of this include moving away, divorce, death, or birth in the family which changes how the family operates.
@@
Centeroid Comment: “Love how OP tries to frame it as "looking out for his son" when in reality, he selfishly wants his son to be a doormat & do what he wants versus being happy. I’m betting the daughter got the same treatment & that’s why she went NC to avoid all his stupid drama. … They’ve grown up & learned to be adults, about damn time you do the same before you torpedo any chances of ever having them in your lives again. …”
@@
Random Comment: “Husband has been clear about not wanting another baby but OP has insisted on carrying to term. I think he’s been clear about not wanting to parent.”
D.5 C5: Relationship Issues with Men
@@
The Relationship Issues with Men cluster contains comments which discuss situations in which people have problems with men.
@@
Centeroid Comment: “I think you and him need to have a serious conversation about this. That behavior is unacceptable.”
@@
Random Comment: “… In short, you should be able to watch your CNN at his house and he should be able to watch his Fox News. This has nothing to do with his cancer diagnosis, nor does it have to do specifically with your relationship to him. …”
D.6 C6: Relationship Issues with Women
@@
The Relationship Issues with Women cluster contains comments which discuss situations in which people have problems with women.
@@
Centeroid Comment: “NTA that sounds a little mean to be honest. I think it’s understandable to be upset by it, but hopefully she will take it seriously and stop doing that.”
@@
Random Comment: “… I would NEVER talk to my boyfriend that way for trying to help me. You were right to leave. You were being an excellent boyfriend, but she was being a jerk.”
D.7 C7: Sympathy & Support